Loading in 2 Seconds...
Loading in 2 Seconds...
Pulsar System Overview and Commissioning Plan Cheng-Ju Lin Fermilab Installation Readiness Review 02/08/2005. Outline: - Items from last review - Pulsar baseline system - Commissioning strategy - From now till March.
Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.
Installation Readiness Review
- Items from last review
- Pulsar baseline system
- Commissioning strategy
- From now till March
of action items that are essential for the commissioning.
IRR committee expressed concerns about the long list of items.
PC,TL2D,RC software issues:1.) Update trigger database to include new Pulsar software tags. (Daniel+database folks: Nov 12)2.) Test new trigger table built script. (Daniel+Kristian: Nov 12)3.) Generate TL2D in PC. (Kristian: Nov 1)4.) Porting command and control code to AMD machine. (Kristian: Oct 22)5.) Order two AMD machines. (JJ: target arrival date Nov 30)6.) Update/test FER code to readout Pulsar TL2D bank. (Cheng-Ju+Jane: Oct 29)7.) Measure readout time for TL2D and other configurations. (Cheng-Ju: Nov 15)8.) Update/test FER code to send scaler info to ScalerMon. (Cheng-Ju+Daniel+DAQ experts: Nov 19)9.) Update electron trigger code. (Kristian: Oct 25)10.) Update RC to switch between Pulsar and Alpha decision, confirm L2 tagset, L2 alpha auto reject. (Badgett, Tom Wright: 1st item is done, the rest complete by Nov 19)Firmware Developments:1.) Isolation firmware improvement. (Sakari: testing ongoing, complete by Nov 5)2.) Implement firmware for additional system timing measurements and source ID. (Sakari: Nov 5)2.) ShowerMax firmware improvement and zero-suppressing. (Sakari: firmware testing ongoing)3.) SLINK package sending improvement to reduce transfer latency. The improvement is being tested on the SVT path first. (Sakari: testing in test-stand is done; Shawn+Sakari: ongoing testing in trigger room)4.) SLINK merger firmware improvement for timing optimization. (Sakari: Nov 10)5.) FREEZE all firmware development on Nov 15. Only bug fixes will be implemented.
ShowerMax:1.) Modify PulsarMon to accommodate new zero-suppression firmware and integrate into official PulsarMon. (Chris: by Oct 29)2.) Finish testing new firmware with ShowerMax "glitch" protection in test-stand and trigger room. (Chris+Sakari: Completed while I was making the list :-)3.) Checkout zero-suppression firmware in trigger room. (Sakari+Chris: Oct 29)ShowerMax Active Splitter:1.) Test partial active splitter in L2 torture with low threshold; understand latent error rate. (Chris: Oct 22)2.) Modify AS to match light output of the splitter with the current system. Test modified boards (VLTN testing). (Chris: Oct 25-29)3.) If 1-2 are okay, install all AS cards/jumper fibers and test complete (splitted) system using patterns, low threshold runs in L2 torture. (Chris: Nov 5-15)PulsarMon:1.) PulsarMon will be integrated into the official TrigMon release on Oct 25. (when Charles Plager returns from his west coast trip) (CP+Kaori: Oct 29)2.) Continue running a parallel TrigMon online (by Pulsar folks) to validate PulsarMon. (Wojtek+Vadim: ongoing)Cluster Board:1.) Continue to work with Monica to resolve upstream protocol violation from CLIQUE board. The CLIQUE schematics were resurrected. Will study it with Monica.2.) Compile low threshold trigger into the new trigger tables (to try to re-produce cluster errors in the trigger room). (Vadim+JDL: Nov 1)3.) Continue to debug Iso-Cluster firmware: - Problem reproducible in L2 torture - New firmware available for testing (Sakari+Vadim+Wojtek: Nov 25)Other Items:1.) Clean up cables in Pulsar and alpha crates. (Cheng-Ju+JDL: Oct 29)2.) The plan is to start re-integrate the Pulsar crate with all paths the week of Nov 15th.
ALL ITEMS ARE DONE!!!
L1 Et Sums
HRR, CONFIG, ETC…
A very talented and hardworking group of people!!!
Red = students
Philosophy: do our best to minimize downtime to CDF physics
Two important items:
- Split all input signal to the global Level-2 crates
(LVDS, HotLink and TAXI fibers). See Chris’ talk
for ShowerMax fiber splitting.
- Have extensive validation and
monitoring tools in place to
ease the task of debugging
the system PulsarMON.
Wojtek with the LVDS splitter board
I. Purely parasitic running:
- We’ve been running in this mode for the most part since Dec. 04.
- The only operational impact is writing data to the event stream.
- Able to debug essentially all firmware and trigger algorithm
bugs in this mode.
- Also useful in the showerMax fiber splitting effort.
II. Semi-parasitic (“symbiotic”) running:
- Allowing Pulsar to handshake with Trigger Supervisor
in physics runs. However, always send L2Reject.
- Pulsar running in “operational-like” environment
- We have been running in this mode fairly smoothly since Feb 2.
Only a few operational glitches were discovered along the way.
- Ran through > billions of L1A.
Some questions from the committee:
Q: What is the pager rotation plan?
We’ll follow the current system and rotate on a weekly basis.
Initially, we’ll also have a local backup pager carrier.
The duty of the pager carrier is limited to the Pulsar system.
However, historically L2 pager carrier also responds to L2
CAL trigger problems (since Monica is offsite).
Q: For how long do we need to maintain the legacy system?
We would prefer that the legacy system be maintained as
hot spare (powered with full support) till at least the summer
shutdown. Afterward, it could be maintained as “cold spare”.
Q: What hardware configuration will be used for operations.
In particular, what signals will be split/fanned-out?
We would argue to maintain the full signal split/fannout
infrastructure until the legacy system becomes cold spare.
Depending on the R&D need, we may want to support the
splitter for an even longer term.
Q: What is your VME crate/test-stand need?
B0l2pu00 will be the primary crate for operation. We’ll need
at least another crate in the trigger room for L2 and XFT R&D.
Right now we have 2 test-stand crates in B0. We would like
to maintain those two crates for R&D work in the immediate